Consider the three stocks in the following table. Pt represents the price at the end of period t and Qt is the number of shares outstanding. Stock C splits 3:1 during period 2.

P0 Q0 P1 Q1 P2 Q2
A $          16.00 150.00 $         18.00 150.00 $   21.00 150.00
B $          32.00 400.00 $         40.00 400.00 $   42.00 400.00
C $          50.00 200.00 $         63.00 200.00 $   29.00 600.00

Calculate the price-weighted index return for the period ending in 2. Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.
Answer #1

IV1 = [P1i] / n

= [$18 + $40 + $63] / 3 = $121 / 3 = $40.33

IV2 = [P2i] / n

= [$21 + $42 + $29] / 3 = $92 / 3 = $30.67

R(year 2) = [IV2 / IV1] - 1

= [$30.67 / $40.33] - 1 = 0.7603 - 1 = -0.2397, or -23.97%